OpenStudy (anonymous):

Simplify the expression. 16x^2-20x-6/4(2x-3) 2(2x + 1) 2x - 3 2x + 1/2 4x + 1/2

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

16x^2-20x-6 tke 2 out 2(8x^2-10x-3)=2(8x^2-12x+2x-3)=2[4x(2x-3)+1(2x-3)]=2(2x-3)(4x+1) so 16x^2-20x-6/4(2x-3) =[2*(2x-3)*(4x+1)]/[4(2x-3)]=(4x+1)/2

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

it looks exactly like https://media.glynlyon.com/g_alg02_ccss_2013/6/img_examb_33.gif

OpenStudy (anonymous):

Thank you

OpenStudy (blank ):

I tried to answer your question but Openstudy stopped working for me. Sorry about that. I got the same as @sidsiddhartha